OpenStudy (anonymous):

Part 1: Explain, in complete sentences, how you would use the substitution method to solve the following system of equations. (3 points) Part 2: Provide the solution to the system. (1 point) 2x – y = 5 x + 2y = 15

OpenStudy (cwrw238):

first rearrange first equation to get y in terms of x: 2x = y + 5 y = 2x - 5 Now substitute y = 2x-5 in second equation: x + 2(2x-5) = 15 x + 4x - 10 = 15 5x = 15+10 = 25 x = 5 Now plug x = 5 into the first equation to find value of y: 2(5) - y = 5 y = 10-5 = 5 solution set is x=5,y=5

OpenStudy (anonymous):

Thanks, i appreciate your help